find a coordinate of a point that lies on the y=3/2x+6
Answer by stanbon(75887)

find a coordinate of a point that lies on the y=(3/2)x+6
---
Let x = 2, then y = (3/2)2+6 = 3+6 = 9
---
Point: (2,9)
===============
Cheers,
Stan H.